How many distinct permutations are there in each word?

20. MOMMY

21. REARRANGE

22. MATHEMATICS

23. PROBABILITY

24. STATISTICS

25. PHILIPPINES


1
Expert's answer
2022-01-19T08:40:58-0500

20) "N=\frac{5!}{3!}=20."

21) "N=\frac{9!}{3!2!2!}=15120."

22) "N=\frac{11!}{2!2!2!}=4989600."

23) "N=\frac{11!}{2!2!}=9979200."

24) "N=\frac{10!}{3!3!2!}=50400."

25) "N=\frac{11!}{3!3!}=1108800."